find the slope, a point on the line, and graph the following equation
y-1=2(x-3)
This is what I did. I want to check my work and ensure I am following directions correctly.
y=2x-5
m=2/1
b=-5
Now I graph a point on the y axis at negative 5 and then rise 2 and run 1, correct?

Answer by Alan3354(69443)

find the slope, a point on the line, and graph the following equation
y-1=2(x-3)
This is what I did. I want to check my work and ensure I am following directions correctly.
y=2x-5
m=2/1
b=-5
Now I graph a point on the y axis at negative 5 and then rise 2 and run 1, correct?
==============
It's correct.
You can pick any value for x and find y.
eg, x = 3, y = 1 --> the point (3,1)